Reaction DataReactants Products
Al(NO3)3 NaCl NaNO3 AlCl3Starting Amount in Reaction 4 moles 9 moles ?
Determine the maximum amount of NaNO3 that was produced during the experiment. Explain how you determined this amount.

1
Expert's answer
2021-02-28T06:22:17-0500

equalize the coefficients:

Al(NO3)3+3NaCl=3NaNO3+AlCl3 

according to the equation 1 mol Al(NO3)3 reacted with 3 moles NaCl. hence, 4 moles Al(NO3)3 should react with 12 moles NaCl, and 9 moles reacted NaCl. So NaCl is limiting, we decide further on this reagent.

"\frac{9}{3}=\frac{x}{3}"

x=9

9 moles NaNO3